# There Was Never a Bottleneck in Concept Bottleneck Models

### v2

@@ 5 Other Fundamental Theoretical Flaws of CBMs
One-vs-Rest strategies exhibit several limitations: (i) individual binary classifiers tend to be biased toward the negative class, and (ii) the predicted probabilities across classifiers are typically uncalibrated [43]. To illustrate these issues, we design an experiment where the concepts are defined based on a four-class spiral dataset with imbalanced class distributions. We train: (i) a CBM with One-vs-Rest binarization, and (ii) an MCBM modeling concepts directly as multiclass variables.
@@ 6 Conclusions
Acknowledgments and Disclosure of Funding
This work has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101007666, MCIN/AEI/10.13039/501100011033 under Grant PID2021-126061OB-C44, and the Government of Aragón (Grant Group T36 23R).
@@ Appendix D Hard Concept Bottleneck Models
As discussed in Section 3, one of the most successful approaches to mitigating information leakage is the family of Hard Concept Bottleneck Models (HCBMs) [18]. Unlike MCBMs, which enforce an Information Bottleneck (IB) directly at the representation level $\bm{z}$, HCBMs operate by pruning the information contained in $\bm{z}$ before producing the task prediction $\hat{\bm{y}}$. In doing so, they encourage $\hat{\bm{y}}$ to depend solely on $\bm{c}$, thereby yielding interventions that are more reliable in practice. This approach differs fundamentally from MCBMs: whereas Equation 4 in MCBMs minimizes over $Z_{j}$, thereby removing nuisances directly from the representation $z_{j}$, HCBMs optimize at the prediction level, focusing on $\hat{Y}$. Formally, they are trained to minimize:
@@
Figure 9: Graphical models of HCBMs with two concepts and two-dimensional representations. Hard Concept Bottleneck Models obtain a binarized version $\hat{c}^{b}_{j}$ of each predicted concept $\hat{c}_{j}$ through the binarizing head $\color[rgb]{1,0,1}p(\hat{c}^{b}_{j}\mid\hat{c}j)$ (fuchsia arrows). Unlike the models in Figure 2, HCBMs predict the task output $\hat{\bm{y}}$ from the binarized concepts $\hat{c}^{b}_{j}$ using the new task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}\mid\bm{z})$ (green arrows).
Appendix E Experiments Details
E.1 Hyperparameters for Section 4
@@
MPI3D Shapes3D CIFAR-10 CUB
$f_{\theta}$ architecture ResNet20 ResNet20 2 conv. layers InceptionV3
$f_{\theta}$ pretraining None None None ImageNet
$g^{y}_{\phi}$ hidden layers 64 64 64 256
low $\gamma$ 1 1 0.1 0.05
medium $\gamma$ 3 3 0.3 0.1
high $\gamma$ 5 5 0.5 0.3
number of epochs 50 50 200 250
batch size 128 128 128 128
optimizer SGD SGD Adam SGD
learning rate $6\times 10^{-3}$ $6\times 10^{-3}$ $1\times 10^{-4}$ $2\times 10^{-2}$
momentum 0.9 0.9 0. 0.9
weight decay $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$
scheduler Step Step Step Step
step size (epochs) 20 20 80 100
scheduler $\gamma$ 0.1 0.1 0.1 0.1
E.2 Datasets
@@
CIFAR-10 is a widely used image classification benchmark consisting of 60,000 natural images of size 32 $\times$ 32, divided into 10 classes (e.g., airplanes, automobiles, birds, cats, etc.). The dataset is split into 50,000 training and 10,000 test images, with balanced class distributions. To reduce the need for manual concept annotations, the concepts are synthetically derived following the methodology of [23]. A total of 143 attributes are extracted using GPT-3 [33]; 64 form the concept set $\bm{c}$, while the rest define the nuisance set $\bm{n}_{y}$. Binary values are obtained with the CLIP model [44] by comparing the similarity of each image to the embedding of an attribute and to its negative counterpart.
@@
The Caltech-UCSD Birds (CUB) dataset contains 11,788 images of 200 bird species, annotated with part locations, bounding boxes, and 312 binary attributes. Following the approach of Koh et al. [9], we retain only the attributes that are present in at least 10 species (based on majority voting), resulting in a filtered set of 112 attributes. These attributes are grouped into 27 semantic clusters, where each group is defined by a common prefix in the attribute names. In our setup, the task variable $\bm{y}$ is to classify the bird species. The concept set $\bm{c}$ consists of the attributes belonging to 12 randomly selected groups (per run), while the nuisance set $\bm{n}y$ includes the attributes from the remaining 20 groups. Since most attributes exhibit some correlation with the classification task, we set $\bm{n}_{\bar{y}}$ to the empty set.

### v3

@@ 5 Other Fundamental Theoretical Flaws of CBMs
One-vs-Rest strategies exhibit several limitations: (i) individual binary classifiers tend to be biased toward the negative class, and (ii) the predicted probabilities across classifiers are typically uncalibrated (Bishop, 2006). To illustrate these issues, we design an experiment where the concepts are defined based on a four-class spiral dataset with imbalanced class distributions. We train: (i) a CBM with One-vs-Rest binarization, and (ii) an MCBM modeling concepts directly as multiclass variables.
@@ 6 Conclusions, Limitations and Future Directions
Regarding the limitations of this work, we highlight the following: (i) MCBMs introduce a new hyperparameter, $\gamma$, which must be tuned to balance predictive accuracy and interpretability; (ii) the representation head $g_{\phi}^{z}$ adds a small number of parameters, though this overhead is negligible relative to the backbone; and (iii) while MCBMs consistently reduce nuisance information, complete removal remains challenging for high-variance datasets, suggesting that performance depends on the expressive capacity of the backbone architecture.
As for future directions, promising avenues include: (i) extending the model with an auxiliary latent variable $z_{m+1}$ appended to $\bm{z}$ to explicitly capture task-relevant nuisance factors $n_{y}$, allowing $z_{1},\dots,z_{m}$ to remain strictly interpretable while still retaining task-useful information in the full representation $\bm{z}$; and (ii) studying how the choice of prior distributions $q_{\phi}(z_{j}\mid c_{j})$ affects representation quality and evaluation metrics such as disentanglement, alignment, and concept leakage.
Acknowledgments
This work has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101007666, MCIN/AEI/10.13039/501100011033 under Grant PID2024-155948OB-C53, and the Government of Aragón (Grant Group T36 23R).
Code Availability
An implementation accompanying this work is available at https://github.com/antonioalmudevar/minimal_cbm
@@ Appendix D Hard Concept Bottleneck Models
As discussed in Section 3, one of the most successful approaches to mitigating information leakage is the family of Hard Concept Bottleneck Models (HCBMs) (Havasi et al., 2022). Unlike MCBMs, which enforce an Information Bottleneck (IB) directly at the representation level $\bm{z}$, HCBMs operate by pruning the information contained in $\bm{z}$ before producing the task prediction $\hat{\bm{y}}$. In doing so, they encourage $\hat{\bm{y}}$ to depend solely on $\bm{c}$, thereby yielding interventions that are more reliable in practice. This approach differs fundamentally from MCBMs: whereas Equation 4 in MCBMs minimizes over $Z_{j}$, thereby removing nuisances directly from the representation $z_{j}$, HCBMs optimize at the prediction level, focusing on $\hat{Y}$. Formally, they are trained to minimize:
@@
Figure 9: Graphical models of HCBMs with two concepts and two-dimensional representations. Hard Concept Bottleneck Models obtain a binarized version $\hat{c}^{b}_{j}$ of each predicted concept $\hat{c}_{j}$ through the binarizing head $\color[rgb]{1,0,1}p(\hat{c}^{b}_{j}\mid\hat{c}_{j})$ (fuchsia arrows). Unlike the models in Figure 2, HCBMs predict the task output $\hat{\bm{y}}$ from the binarized concepts $\hat{c}^{b}_{j}$ using the new task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}\mid\hat{\bm{c}}^{b})$ (green arrows).
Appendix E Other Intervention Procedures
As detailed in Section 4.3, we further examine how performance evolves as we intervene on an increasing number of randomly selected concepts, with the resulting curves shown in Figure 10. The overall behavior is consistent with the trends observed in Figure 5, highlighting similar model characteristics. As expected, the results exhibit somewhat higher variance across seeds, since each run intervenes on a different subset of concepts.
(a) CIFAR-10
(b) CUB
(c) AwA2
Figure 10: Error (y-axis) versus percentage of concepts intervened (x-axis) across different models for randomly selected concepts.
Appendix F Experiments Details
F.1 Hyperparameters for Section 4
@@
MPI3D Shapes3D CIFAR-10 CUB AwA2
$f_{\theta}$ architecture ResNet20 ResNet20 2 conv. layers InceptionV3 ResNet-50
$f_{\theta}$ pretraining None None None ImageNet ImageNet
$g^{y}_{\phi}$ hidden layers 64 64 64 256 256
$g^{c}_{\phi}$ hidden layers None None None None None
$g^{z}_{\phi}$ hidden layers [^1] Beyond the choice of $\gamma$, MCBMs introduce an additional design decision: the architecture of $g_{\phi}^{z}(c_{j})$. In all our experiments, we implement this module as a small MLP with a single hidden layer of size 3, adding only 8 parameters per concept. Since concept sets typically contain at most 200 concepts, this corresponds to roughly 1600 additional parameters—negligible compared to the size of standard neural encoders $f_{\theta}$. 3 3 3 3 3
low $\gamma$ 1 1 0.1 0.05 0.05
medium $\gamma$ 3 3 0.3 0.1 0.1
high $\gamma$ 5 5 0.5 0.3 0.3
number of epochs 50 50 200 250 120
batch size 128 128 128 128 128
optimizer SGD SGD Adam SGD SGD
learning rate $6\times 10^{-3}$ $6\times 10^{-3}$ $1\times 10^{-4}$ $2\times 10^{-2}$ $2\times 10^{-2}$
momentum 0.9 0.9 0. 0.9 0.9
weight decay $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$ $4\times 10^{-5}$
scheduler Step Step Step Step Step
step size (epochs) 20 20 80 100 50
scheduler $\gamma$ 0.1 0.1 0.1 0.1 0.1
F.2 Datasets
@@
CIFAR-10 is a widely used image classification benchmark consisting of 60,000 natural images of size 32 $\times$ 32, divided into 10 classes (e.g., airplanes, automobiles, birds, cats, etc.). The dataset is split into 50,000 training and 10,000 test images, with balanced class distributions. To reduce the need for manual concept annotations, the concepts are synthetically derived following the methodology of (Vandenhirtz et al., 2024). A total of 143 attributes are extracted using GPT-3 (Brown et al., 2020); 64 form the concept set $\bm{c}$, while the rest define the nuisance set $\bm{n}_{y}$. Binary values are obtained with the CLIP model (Radford et al., 2021) by comparing the similarity of each image to the embedding of an attribute and to its negative counterpart.
@@
The Caltech-UCSD Birds (CUB) dataset contains 11,788 images of 200 bird species, annotated with part locations, bounding boxes, and 312 binary attributes. Following the approach of Koh et al. (2020), we retain only the attributes that are present in at least 10 species (based on majority voting), resulting in a filtered set of 112 attributes. These attributes are grouped into 27 semantic clusters, where each group is defined by a common prefix in the attribute names. In our setup, the task variable $\bm{y}$ is to classify the bird species. The concept set $\bm{c}$ consists of the attributes belonging to 12 randomly selected groups (per run), while the nuisance set $\bm{n}_{y}$ includes the attributes from the remaining 15 groups. Since most attributes exhibit some correlation with the classification task, we set $\bm{n}_{\bar{y}}$ to the empty set.
AwA2
The Animals with Attributes 2 (AwA2) dataset (Xian et al., 2017) contains 37,322 images of 50 animal classes annotated with 85 human-defined attributes describing appearance, behavior, and habitat. In our setup, the task variable $\bm{y}$ is the animal class. To construct the concept and nuisance partitions, we retain a subset of 20 attributes—covering fundamental appearance and morphology features—as the concept set $\bm{c}$, while the remaining 65 attributes form the nuisance set $\bm{n}_{y}$. Since nearly all attributes exhibit some degree of correlation with the class label, we set $\bm{n}_{\bar{y}}$ to the empty set. Following the preprocessing in Xian et al. (2017), we binarize continuous attributes using a threshold of 0.5.